# Less is More: Improving LLM Alignment via Preference Data Selection

### v2

@@ Abstract
Direct Preference Optimization (DPO) has emerged as a promising approach for aligning large language models with human preferences. While prior work mainly extends DPO from the aspect of the objective function, we instead improve DPO from the largely overlooked but critical aspect of data selection. Specifically, we address the issue of parameter shrinkage caused by noisy data by proposing a novel margin-maximization principle for dataset curation in DPO training. To accurately estimate margins for data selection, we propose a dual-margin guided approach that considers both external reward margins and implicit DPO reward margins. Extensive experiments demonstrate that our method reduces computational cost dramatically while improving performance. Remarkably, by using just 10% of the Ultrafeedback dataset, our approach achieves 3% to 8% improvements across various Llama and Mistral series models on the AlpacaEval 2.0 benchmark. Furthermore, our approach seamlessly extends to iterative DPO, yielding a roughly 3% improvement with 25% online data, while further reducing training time. These results highlight the potential of data selection strategies for advancing preference optimization.
Machine Learning, ICML
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF; Christiano et al., 2017; Ziegler et al., 2019) has emerged as a crucial technique for aligning Large Language Models (LLMs) with human preferences and values. Traditional RLHF implementations involve a two-stage process: reward model training based on preference data followed by reinforcement learning optimization. However, this approach presents significant computational challenges, requiring loading multiple model instances and extensive hyperparameter tuning.
As an alternative, Rafailov et al. (2024) introduced Direct Preference Optimization (DPO), which streamlines the alignment process by directly optimizing the LLM policy from preference data. DPO has demonstrated comparable effectiveness while substantially reducing computational requirements compared to classical RLHF. Following DPO’s introduction, numerous studies have proposed improvements through modified learning objectives (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024) and iterative learning schemes (Xiong et al., 2024). While these algorithmic advances have shown promise, there remains a critical gap in our understanding of the data-centric aspects of preference learning: what characteristics of preference data contribute most to model alignment?
@@
(1) We prove in theory the necessity of data selection in the presence of exogenous noise. Specifically, the inherent noise in the reward model may flip the preference between response pairs, leading to the emergence of the parameter shrinkage issue. Furthermore, we demonstrate that margin-based selection criteria can effectively address this issue by inducing parameter inflation.
(2) Building on this insight, we propose a margin-maximization principle for dataset curation that incorporates signals from the ensemble of external rewards and DPO implicit rewards. Through extensive experiments across diverse datasets and base models, we show that this selection strategy shows two consistent advantages: it substantially reduces computational overhead via efficient data selection and improves model performance compared to training on the full dataset. In particular, on the Ultrafeedback dataset and its variants, our method identifies a $10\%$ data subset for DPO training on LLama and Mistral series models, consistently achieving 3% to 8% point improvements on the AlpacaEval 2.0 benchmark relative to training on the complete dataset.
@@ 2 Background
Reward Learning. In the reward learning stage, a reward model is trained to approximate human preferences based on preference data. By adopting the Bradley-Terry model (Bradley & Terry, 1952) to capture human preference, reward training involves minimizing the loss:
@@
Policy Optimization with Reinforcement Learning. Once the reward model $r$ is trained, it is used to guide the optimization of a policy $\pi_{\theta}(y|x)$, where $\theta$ denotes the parameters of the model. This stage often employs reinforcement learning techniques such as Proximal Policy Optimization (PPO; Schulman et al., 2017) to optimize the policy by maximizing the expected reward.
@@
Recently, as an alternative to the RL-based policy optimization in RLHF, Direct Preference Optimization (DPO; Rafailov et al., 2024) has been proposed. DPO simplifies the reward alignment process by directly incorporating human preference data into supervised training. Instead of defining and optimizing a reward function explicitly, DPO minimizes
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\mathrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\mathrm{ref}}(y_{l}|x)}\right).$
@@ 3 Methodology
We follow the model from Zhu et al. (2023) to illustrate why data selection can improve model performance. We assume that reward model $r(x,y)=\langle\phi(x,y),\omega^{*}\rangle$ with some feature function $\phi(\cdot,\cdot)$. For reward learning, our reward model can be an explicit $r(x,y)$ (Ouyang et al., 2022), while for DPO, $\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}$ plays the role of reward model implicitly (Rafailov et al., 2024). Based on observations in previous literature, we can derive such features by removing the last layer of the pre-trained model. However, both humans and other large models may use inaccurate reward functions to generate labels, where the labels represent the ranking of two responses. We say the preference between $y_{w}$ and $y_{l}$ is generated by $r(x,y_{w})-r(x,y_{l})+\zeta$ where $\zeta$ is an exogenous error. We use $\Delta\phi(x)$ to denote $\phi(x,y_{w})-\phi(x,y_{l})$ for simplicity.
@@
$\mathcal{L}_{\mathrm{RM}}(\omega)$ $=-\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\log(\frac{1}{1+e^{-\langle\Delta\phi(x),\omega\rangle}})$
$+\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\log(\frac{1}{1+e^{\langle\Delta\phi(x),\omega\rangle}})\Big{]}.$ (1)
@@
$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}$
$=$ $\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$ (2)
@@
$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}$
$\geq$ $\mathbb{E}_{x}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$
@@
$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}$
$\leq$ $\mathbb{E}_{x}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$
@@
We provide the underlying intuition with an extreme example. If $\mathbb{V}(\zeta)$ goes to infinity, the preference between $y_{w}$ and $y_{l}$ mainly depends on $\zeta$, approaching a Rademacher distribution, then $\omega=0$ could be a good solution to Section 3. In other words, $\zeta$ offsets part of the information provided by the reward model, causing the model’s parameters to shrink toward zero. Thus, data selection is essential for acquiring policies with good performance. Finally, we remark that $\zeta$ can come from multiple resources, including human classification errors, different embeddings or reward models from other LLMs and so on.
@@
First, when the margin is large, namely, $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$ is large, from the S-shaped graph of $\sigma(\cdot)$, we know that the slope is very small in this area. As a result, the probability of preference reversal caused by $\zeta$ is low, which means the likelihood of incorrect samples is also low.
Secondly, given prompt $x$, as we select data with large $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$, the posterior distribution of $\zeta$ is skewed toward the positive side. Therefore, the preferences corresponding to this kind of data are more pronounced, leading to inflated estimates of $\omega$ in Section 3.
Finally, we point out that if realized $y_{w}$ and $y_{l}$ are all separable, proportional scaling of $\omega$ can reduce the value of Section 3 continuously. Hence, some techniques like regularization or early stopping when training are indispensable.
@@
In summary, inaccuracies in the reward model can cause the parameters of LLMs to shrink toward zero. By selecting data with larger margins, we can compensate for the performance degradation caused by this shrinkage. The balance between parameter shrinkage and inflation offers the potential to enhance the performance of LLMs.
3.1 Dual-Margin guided Data Selection for Preference Learning
In this section, we present the Dual-Margin (DM) guided efficient preference data selection that enhances the performance of direct preference optimization while significantly reducing the training cost. Driven by the theoretical result, our main idea is to let the model reach an overall high margin during preference learning. We realize this by providing a robust estimation of reward margins for the entire dataset ahead of full-set training, which then allows efficient high-margin data filtering. DM operates in two stages: (1) computing external and implicit rewards, and (2) fusing these rewards to suppress noise and achieve more reliable margin estimates.
Initialization.
We aim select a subset $\mathcal{D}_{\mathrm{train}}$ from $\mathcal{D}=\{(x^{i},y_{w}^{i},y_{l}^{i})\}_{i=1}^{N}$, such that the model $\pi_{\theta}$ trained on the $\mathcal{D}_{\mathrm{train}}$ achieves performance comparable or better than that trained on $\mathcal{D}$. To guide this selection, we leverage an external reward model $r_{\mathrm{ex}}(x,y)$ and a small seed dataset $\mathcal{D}_{0}$ to learn an implicit reward model $r_{\mathrm{im}}(x,y)$.
Step 1: Reward calculation.
First, we calculate $r_{\mathrm{ex}}(x^{i},y_{w}^{i})$ and $r_{\mathrm{ex}}(x^{i},y_{l}^{i})$ for each datum in $\mathcal{D}$. Rafailov et al. (2024) proves that we can calculate implicit reward as
$r_{\mathrm{im}}(x,y)=\log\frac{\pi_{\theta}(y|x)}{{\pi_{\mathrm{ref}}}(y|x)}=\log\frac{\prod_{j}\pi_{\theta}(y_{j}|x,y_{1,\cdot,j-1})}{\prod_{j}\pi_{\mathrm{ref}}(y_{j}|x,y_{1,\cdot,j-1})}.$
As this reward margin between chosen and rejected samples calculated by different models shows a strong correlation (refer to Figure LABEL:fig:visual), we propose disentangling its calculation from the target model $\pi_{\theta}$. In particular, we introduce a small model $\pi_{s}$, and fine-tune it on $\mathcal{D}_{0}$ to get a weakly aligned model $\pi_{s}^{1}$. These two models are then utilized for low-cost implicit rewards calculation on $\mathcal{D}$.
Step 2: Margin fusion.
After calculating the margins $m_{\mathrm{ex}}=r_{\mathrm{ex}}(x,y_{w})-r_{\mathrm{ex}}(x,y_{l})$ and $m_{\mathrm{im}}=r_{\mathrm{im}}(x,y_{w})-r_{\mathrm{im}}(x,y_{l})$, we consider appropriate fusion strategies for combining these two signal sources to reduce noise in the reward modeling process. The simplest approach, which we call DM-ADD, directly sums the signals. This strategy reflects a lenient selection criterion where samples are considered valuable if they demonstrate a high margin from either reward source.
A strict fusion approach considers that samples should receive low priority if they show a low margin in either signal. We first transform margin values to margin-guided probability through a simple linear transformation
$\mathbb{P}(m)=\frac{\mathrm{clip}(m,M_{1},M_{2})-M_{1}}{M_{2}-M_{1}},\quad\text{ for }m\in\{m_{\mathrm{ex}},m_{\mathrm{im}}\},$
where $\mathrm{clip}(m)=\min(\max(m,M_{1}),M_{2}))$ and $(M_{1},M_{2})$ are tuning parameters. Following the fusion principle in multi-view probabilistic clustering (Deng et al., 2024), we further obtain
$\mathbb{P}(y_{w}\geq y_{l}|m_{\mathrm{ex}},m_{\mathrm{im}})$
$=$ $\frac{\mathbb{P}(m_{\mathrm{ex}})\mathbb{P}(m_{\mathrm{im}})}{\mathbb{P}(m_{\mathrm{ex}})\mathbb{P}(m_{\mathrm{im}})+(1-\mathbb{P}(m_{\mathrm{ex}}))\cdot(1-\mathbb{P}(m_{\mathrm{im}}))}.$ (3)
This adaptive approach mitigates the adverse effects of outlier samples with unusually high margin values. See more implementation details in Appendix B. Overall, this fusion approach is referred to as DM-MUL.
Sample Selection.
We select the samples with the highest fused margins to construct $\mathcal{D}_{\mathrm{train}}$. After selection, the subset can be used for DPO training on any target model.
3.2 Data Efficient Online Iterative DPO
Online RLHF (Xiong et al., 2024) has gained significant attention for its effectiveness in aligning models with human preferences. The method iteratively generates multiple responses for given prompts and employs an external reward model to identify Best-of-N (BoN) and Worst-of-N (WoN) responses, forming preference pairs. This ‘self-reward maximization’ approach effectively sharpens the model’s response space by reducing the probability of generating low-quality responses (Huang et al., 2024). However, the online generation process still produces ambiguous samples, particularly when prompts either have highly deterministic answer patterns or are exceptionally challenging for models to address. Such on-policy data pairs may not effectively guide model improvement. To address this problem, we propose a straightforward solution called DPO-DM: incorporating DM into each round of the on-policy data collection process and conducting preference optimization exclusively on selected high-margin samples. This approach can substantially reduce training costs, particularly when dealing with large prompt spaces.

### v3

@@ Abstract
Direct Preference Optimization (DPO) has emerged as a promising approach for aligning large language models with human preferences. While prior work mainly extends DPO from the aspect of the objective function, we instead improve DPO from the largely overlooked but critical aspect of data selection. Specifically, we address the issue of parameter shrinkage caused by noisy data by proposing a novel margin-maximization principle for dataset curation in DPO training. To further mitigate the noise in different reward models, we propose a Bayesian Aggregation approach that unifies multiple margin sources (external and implicit) into a single preference probability. Extensive experiments in diverse settings demonstrate the consistently high data efficiency of our approach. Remarkably, by using just 10% of the Ultrafeedback dataset, our approach achieves 3% to 8% improvements across various Llama, Mistral, and Qwen models on the AlpacaEval2 benchmark. Furthermore, our approach seamlessly extends to iterative DPO, yielding a roughly 3% improvement with 25% online data, revealing the high redundancy in this presumed high-quality data construction manner. These results highlight the potential of data selection strategies for advancing preference optimization.
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF; christiano2017deep, ; ziegler2019fine, ) has emerged as a crucial technique for aligning Large Language Models (LLMs) with human preferences and values. Traditional RLHF implementations involve a two-stage process: reward model training based on preference data followed by reinforcement learning optimization. However, this approach presents significant computational challenges, requiring loading multiple model instances and extensive hyperparameter tuning.
As an alternative, (rafailov2024direct, ) introduced Direct Preference Optimization (DPO), which streamlines the alignment process by directly optimizing the LLM policy from preference data. DPO has demonstrated comparable effectiveness while substantially reducing computational requirements compared to classical RLHF. Following DPO’s introduction, numerous studies have proposed improvements through modified learning objectives (zhao2023slic, ; azar2024general, ; ethayarajh2024kto, ) and iterative learning schemes (xiong2024iterative, ). While these algorithmic advances have shown promise, there remains a critical gap in our understanding of the data-centric aspects of preference learning: what characteristics of preference data contribute most to model alignment?
@@
(1) We prove in theory the necessity of data selection in the presence of exogenous noise. Specifically, the noise in the reward model may flip the preference between response pairs, leading to the emergence of the parameter shrinkage issue. Furthermore, we demonstrate that margin-based selection criteria can effectively address this issue by inducing parameter inflation.
(2) Driven by the theoretical results and the derived margin-maximization principle, we propose a Bayesian Aggregation for Preference data Selection (BeeS) strategy. BeeS incorporates signals from both external rewards and DPO implicit rewards, and deprioritizes a preference pair if it exhibits a low reward margin from any single reward source to mitigate potential noise. Through extensive experiments across diverse datasets and models, we show that this selection strategy shows two consistent advantages: it substantially reduces computational overhead via efficient data selection and improves model performance compared to training on the full dataset. In particular, on the UltraFeedback dataset and its variants, our method identifies a $10\%$ data subset for DPO training on LLama, Mistral, and Qwen series models, consistently achieving 3% to 8% point improvements on the AlpacaEval 2.0 benchmark relative to training on the complete dataset.
@@
1.1 Related Work
Preference learning algorithms. Reinforcement Learning from Human Feedback also known as dueling RL (pacchiano2021dueling, ) or preference-based RL (chen2022human, ), has become a crucial component of recent Large Language Models (LLMs) such as ChatGPT (ouyang2022training, ). While the classical RLHF pipeline traditionally uses Proximal Policy Optimization, several alternative approaches have been proposed. These include but not limited other RL-based training algorithms (li2023remax, ; zhong2024dpo, ), rejection sampling (dong2023raft, ; gulcehre2023reinforced, ), conditional supervised fine-tuning (lu2022quark, ; yang2024rewards, ; zhang2024reward, ), and Direct Preference Optimization (rafailov2024direct, ). Among these alternatives, DPO has gained significant attention due to its simplicity and robust performance. Following the introduction of DPO, numerous works (zhao2023slic, ; azar2024general, ; ethayarajh2024kto, ; meng2024simpo, ; tang2024generalized, ; han2024f, ; xu2024contrastive, ; hong2024orpo, ; wu2024self, ) have attempted to improve its performance by modifying the DPO objective.
Data selection in LLM Fine-tuning. Data selection is crucial in LLM post-training (wang2024survey, ) for two key observations: post-training typically converges rapidly, and excessive data can degrade model performance through overfitting or exposure to toxic content (swayamdipta2020dataset, ; deng2023counterfactual, ). Recent research has focused on enhancing instruction tuning efficiency by identifying high-quality subsets from large instruction datasets (cao2023instruction, ; li2024superfiltering, ; xia2024less, ), often adapting active learning query strategies (ren2021survey, ) to assess sample uncertainty and diversity. However, data efficiency in preference learning remains relatively unexplored. Prior studies have studied reducing annotation costs in preference dataset creation (muldrew2024active, ; yasunaga2024alma, ) and on scenarios involving numerous ranking annotations (thekumparampil2024comparing, ; mukherjee2024optimal, ). Other research aims to improve a model’s ability to distinguish between two responses by adding a margin to the loss term (meng2024simpo, ; park2024disentangling, ; azar2024general, ). Additionally, concurrent work highlights the importance of margins for preference data filtering, though there is debate on whether hard samples help or hinder preference learning (wu2024beta, ; khaki2024rs, ; yu2025rip, ; gao2025principled, ).
Our work firstly provides clear criteria for identifying informative samples while filtering toxic ones, thereby improving both DPO’s efficiency and performance. Furthermore, our method extends to iterative DPO (xiong2024iterative, ) and its variants (zhang2024self, ), wherein training data is dynamically generated by the model during its iterative training process.
@@ 2 Background
Reward Learning. In the reward learning stage, a reward model is trained to approximate human preferences based on preference data. By adopting the Bradley-Terry model (bradley1952rank, ) to capture human preference, reward training involves minimizing the loss:
@@
Policy Optimization with Reinforcement Learning. Once the reward model $r$ is trained, it is used to guide the optimization of a policy $\pi_{\theta}(y|x)$, where $\theta$ denotes the parameters of the model. This stage often employs reinforcement learning techniques such as Proximal Policy Optimization (PPO; schulman2017proximal, ) to optimize the policy by maximizing the expected reward.
@@
Recently, as an alternative to the RL-based policy optimization in RLHF, Direct Preference Optimization (DPO; rafailov2024direct, ) has been proposed. DPO simplifies the reward alignment process by directly incorporating human preference data into supervised training. Instead of defining and optimizing a reward function explicitly, DPO minimizes
$\mathcal{L}_{\mathrm{DPO}}(\theta)=-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\mathrm{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\mathrm{ref}}(y_{l}|x)}\right).$
@@ 3 Methodology
3.1 Parameter Shrinkage and Inflation Analysis
We follow the model from (zhu2023principled, ) to illustrate why data selection can improve model performance. We assume that reward model $r(x,y)=\langle\phi(x,y),\omega^{*}\rangle$ with some feature function $\phi(\cdot,\cdot)$. For reward learning, our reward model can be an explicit $r(x,y)$ (ouyang2022training, ), while for DPO, $\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}$ plays the role of reward model implicitly (rafailov2024direct, ). Based on observations in previous literature, we can derive such features by removing the last layer of the pre-trained model. However, both humans and other large models may use inaccurate reward functions to generate labels, where the labels represent the ranking of two responses. We say the preference between $y_{w}$ and $y_{l}$ is generated by $r(x,y_{w})-r(x,y_{l})+\zeta$ where $\zeta$ is an exogenous error. We use $\Delta\phi(x)$ to denote $\phi(x,y_{w})-\phi(x,y_{l})$ for simplicity.
@@ 3.1 Parameter Shrinkage and Inflation Analysis
$\mathcal{L}_{\mathrm{RM}}(\omega)=-\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\log(\frac{1}{1+e^{-\langle\Delta\phi(x),\omega\rangle}})+\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\log(\frac{1}{1+e^{\langle\Delta\phi(x),\omega\rangle}})\Big{]}.$ (1)
@@
$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}=\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$ (2)
@@
$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle+\zeta}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}\geq\mathbb{E}_{x}\Big{[}\frac{1}{1+e^{\langle\Delta\phi(x),\omega^{*}\rangle}}\frac{e^{\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$
@@
$\mathbb{E}_{x,\zeta}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle-\zeta}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}\leq\mathbb{E}_{x}\Big{[}\frac{1}{1+e^{-\langle\Delta\phi(x),\omega^{*}\rangle}}\frac{e^{-\langle\Delta\phi(x),\omega\rangle}\Delta\phi(x)}{1+e^{-\langle\Delta\phi(x),\omega\rangle}}\Big{]}.$
@@
We provide the underlying intuition with an extreme example. If $\mathbb{V}(\zeta)$ goes to infinity, the preference between $y_{w}$ and $y_{l}$ mainly depends on $\zeta$, approaching a Rademacher distribution, then $\omega=0$ could be a good solution to Equation 1. In other words, $\zeta$ offsets part of the information provided by the reward model, causing the model’s parameters to shrink toward zero. Thus, data selection is essential for acquiring policies with good performance. Finally, we remark that $\zeta$ can come from multiple resources, including human classification errors, different embeddings or reward models from other LLMs and so on.
@@
First, when the margin is large, namely, $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$ is large, from the S-shaped graph of $\sigma(\cdot)$, we know that the slope is very small in this area. As a result, the probability of preference reversal caused by $\zeta$ is low, which means the likelihood of incorrect samples is also low. Secondly, given prompt $x$, as we select data with large $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$, the posterior distribution of $\zeta$ is skewed toward the positive side. Therefore, the preferences corresponding to this kind of data are more pronounced, leading to inflated estimates of $\omega$ in Equation 1. Finally, we point out that if realized $y_{w}$ and $y_{l}$ are all separable, proportional scaling of $\omega$ can reduce the value of Equation 1 continuously. Hence, some techniques like regularization or early stopping when training are indispensable.
@@
In summary, inaccuracies in the reward model can cause the parameters of LLMs to shrink toward zero. By selecting data with larger margins, we can compensate for the performance degradation caused by this shrinkage. The balance between parameter shrinkage and inflation offers the potential to enhance the performance of LLMs. Driven by this theoretical result, our main idea is to let the model reach an overall high margin during preference learning. We realize this by providing a robust estimation of reward margins for the entire dataset ahead of full-set training, which then allows efficient high-margin data filtering.
3.2 Multi-source Margin Aggregation
Building on our previous analysis, we aim to develop a data selection strategy based on the margin-maximization principle, with the calculation of reward margin being the critical component. We examine two distinct types of reward margin calculations: external reward margin and implicit reward margin. The external reward margin is provided by an external reward function, while the implicit reward margin is derived from the implicit reward $\log\frac{\pi_{\theta}}{\pi_{\mathrm{ref}}}$, where $\pi_{\theta}$ represents the policy trained by DPO (rafailov2024direct, ) (see Section 4.1 for details on margin calculation). We visualize the joint distribution of different reward margin sources using the UltraFeedback dataset (cui2023ultrafeedback, ) in Figure LABEL:fig:visual.
  • The left and middle panels of Figure LABEL:fig:visual reveal several key phenomena: (1) The correlation between implicit and external reward margins is notably weak. In particular, samples exhibiting high positive implicit margins span a broad range of external margins, including strongly negative values, and vice versa. This divergence highlights the distinct preference patterns captured by these two reward types. This underscores the need to combine both reward types for a reliable margin estimation. (2) In contrast, we observe a strong correlation between implicit reward margins calculated by models of different sizes (Llama-3.2 3B and 1B). Notably, these two patterns remain consistent across other datasets as well (see Appendix B.2).
  • Online RLHF (xiong2024iterative, ; dong2025self, ) employs the target model to generate multiple responses for given prompts iteratively and uses an external reward model to identify the response pair with the largest margin for DPO training. The right panel of Figure LABEL:fig:visual illustrates that a max-margin pair construction method, even when derived from one strong reward model, can still yield ambiguous preferences when evaluated by another reward model (which shows similar performance on RewardBench Leadboard). This ambiguity in preference signal indicates that the online data generation process can still cause high redundancy, which may offer little to no benefit, or could even be detrimental, to online-DPO training.
These observations highlight the need for multi-source margin aggregation to achieve a more robust margin estimation, thereby enhancing data selection and preference learning. To this end, we propose a strict aggregation strategy, Bayesian Aggregation for Preference data Selection (BeeS), that deprioritizes a preference pair if it exhibits a low reward margin from any single reward source. We implement this method through a general three-step procedure.
Step 1: in-distribution pre-DPO training. Our objective is to obtain the in-distribution implicit reward model with low computational and GPU-memory cost. Given the strong correlation of that margin across different models (See Figure LABEL:fig:visual), this weak-to-strong guidance is feasible. To achieve this, we randomly select a small seed dataset $\mathcal{D}_{0}$ from $\mathcal{D}$ and employ a (or several) small model to perform preference learning on this seed set. The high sample and training efficiency of the DPO loss (kim2025spread, ) ensures the feasibility of this approach.
Step 2: margin calculation. We calculate external and implicit reward margins as $m_{\mathrm{ex}}=r_{\mathrm{ex}}(x^{i},y_{w}^{i})-r_{\mathrm{ex}}(x^{i},y_{l}^{i})$ and $m_{\mathrm{im}}=r_{\mathrm{im}}(x^{i},y_{w}^{i})-r_{\mathrm{im}}(x^{i},y_{l}^{i})$ for each datum in $\mathcal{D}$, where we directly calculate $r_{\mathrm{im}}(x,y)$ as $\log\frac{\pi_{\theta}(y|x)}{{\pi_{\mathrm{ref}}}(y|x)}$. Here, $\pi_{\mathrm{ref}}$ and $\pi_{\theta}$ denote the small model before and after preference learning. We assume that there are $K$ margins involved from these two typical sources.
Step 3: Bayesian aggregation. To mitigate noise from individual reward sources, we utilize Bayesian probability theory for their robust aggregation. Given that reward margins from different sources often vary in their underlying distributions and value ranges, we propose projecting these diverse margins into a unified probability space. This projection serves to quantify the confidence that a specific preference direction $y_{w}>y_{l}$ is correct, formally expressed as $\mathbb{P}(y_{w}>y_{l}|m^{1},m^{2},\cdots,m^{K})$. Assuming that these sources are conditionally independent, then following the Bayesian formula transformation in previous work (liu2022mpc, ; deng2024a3s, ), the preference probability can be expressed as:
$\mathbb{P}(y_{w}>y_{l}|m^{1},m^{2},\cdots,m^{K})=\frac{\prod_{i=1}^{K}\mathbb{P}(y_{w}>y_{l}|m^{i})}{\prod_{i=1}^{K}\mathbb{P}(y_{w}>y_{l}|m^{i})+\prod_{i=1}^{K}(1-\mathbb{P}(y_{w}>y_{l}|m^{i}))}.$ (3)
The typical absence of well-defined preference labels (e.g., $y_{w}>y_{l}$ for clear preference, or $y_{w}=y_{l}$ for indifference) renders the rigorous estimation of single-source preference probabilities challenging. [^1] This estimation typically relies on methods such as isotonic regression or histogram analysis. Consequently, we approximate the probability using a linear projection: $\mathbb{P}(y_{w}>y_{l}|m^{i})=\frac{\mathrm{clip}(m^{i},L,U)-L}{U-L}$ , where $\mathrm{clip}(m,L,U)=\min(\max(m,L),U))$ and $L$, $U$ are tuning parameters. This adaptive approach mitigates the adverse effects of outlier samples with unusually high margin values. See more implementation details and discussion about the derivation approximation of Eq. (3) in Appendix A.1.
Sample selection. We consider the data selection for both the one-pass DPO training and the iterative DPO workflow. For the former, we directly select the samples with the highest aggregated preference probability to construct $\mathcal{D}_{\mathrm{train}}$. The threshold depends on how many preference samples we prefer to use (but should guarantee that samples with negative margins are excluded). For the latter, we only need to train implicit reward models in the first iteration, and BeeS three-step procedures are applied for each iteration.